# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v2

@@ IV-H Impact of Multi-Tenancy
Figure 6c compares the average latency of offloading with on-device processing as the number of co-located applications increases. We observe that when only a single application is deployed on the edge, offloading yields lower latency than on-device processing. However, as the number of co-located applications increases, offloading latency also rises. This increase is caused by resource contention among applications because the A2 GPU does not support isolation. As a result, there is a crossover point at 18 applications, beyond which on-device processing becomes more efficient. Note that the on-device processing latency remains stable regardless of the number of concurrent applications as each device operates independently. In all cases, our models closely predict the latency for both execution strategies.
@@
Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
V Models in Action
@@
In this section, we present a resource manager that leverages our analytic models to adaptively switch between offloading and on-device processing. We describe its decision algorithm and demonstrate its effectiveness through two case studies showcasing how the models enable adaptation to network variability and multi-tenant interference at the edge.
V-A Model-Driven Adaptive Resource Management
The resource manager runs locally on the device and periodically collects runtime metrics, including network bandwidth, edge server load, and request arrival rate. At each epoch, it inputs these metrics into our models to estimate the average latency of both execution strategies and selects the one with the lower predicted latency to execute requests.
Algorithm 1 outlines the decision-making process for selecting the execution strategy. Lines 1-2 predict the on-device average execution latency based on the current arrival rate and the estimated service time on the device. Lines 3-6 compute the predicted average latency for offloading to each edge server, considering server load, service rates, and network conditions. Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
@@ V-A Model-Driven Adaptive Resource Management
// Predict on-device processing latency
@@ 10       $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
V-B Case 1: Fluctuating Network Conditions
@@
Figure 7: Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
We first demonstrate how the resource manager leverages our models to adapt to changing network conditions. We use TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS. Initially, the device has a network bandwidth of 20 Mbps to represent a stable 5G connection. To simulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation, further reduce it to 2 Mbps at t = 40 to emulate severe network congestion or poor coverage, and then restore it to 20 Mbps at t = 60 to represent signal recovery.
@@
Figure 7 shows the observed latency using the execution strategies selected by the resource manager, alongside the model-predicted latencies for both offloading and on-device processing. At t = 0, the predicted offloading latency is lower than on-device processing, therefore the manager chooses to offload requests. At t = 20, despite the reduced bandwidth from 20 to 10 Mbps, the predicted latency for offloading remains lower, and the strategy remains unchanged. However, at t = 40, the drop to 2 Mbps causes the predicted offloading latency (45 ms) to exceed the predicted on-device processing time (21 ms), and the resource manager to switch to on-device processing. Finally, at t = 60, as the network returns to 20 Mbps, the resource manager switches back to offloading. Throughout this sequence, the models closely predict the observed values, validating their accuracy.
@@
Key takeaway. Our models enable the resource manager to dynamically adapt to changing network conditions by switching between execution strategies.
V-C Case 2: Multi-tenant Edge Servers
Figure 8: Observed latency under varying edge server workloads. At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS. At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
Our second case study demonstrates how the resource manager leverages our models to adapt to workload changes across multiple multi-tenant edge servers. In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application. Initially, the request rates of each server are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$. At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50. Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50.
@@ V-C Case 2: Multi-tenant Edge Servers
Figure 8 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing. Initially, from t = 0 to t = 80, the predicted latency for offloading to $E_{1}$ is the lowest, therefore the resource manager offloads requests to $E_{1}$. However, at t = 80, as the workload on $E_{1}$ increases, its predicted latency rises. In response, the resource manager adapts by offloading to the less loaded server $E_{2}$. At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and the resource manager shifts to on-device processing. This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workload.
@@
VI Related Work
Edge offloading. Edge offloading is a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure [5, 8, 43, 44, 45, 6, 46, 47, 48, 49, 50, 51, 52]. For example, Furion [43] offloads computation-intensive components of mobile AR pipelines to the edge. Li et al. [44] introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints. In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators. Our models apply to a wide range of workloads and generalizes to device-edge collaborative processing as shown in the evaluation.
Model-aware placement and resource allocation. Queuing models have been widely used to predict the performance of traditional CPU-bound workloads [53, 3, 54, 55, 56, 57, 20]. More recently, there has been growing interest in modeling the performance of hardware accelerators. For example, Merck et al. [58] model the latency of GPU batching in DNN inference workloads. Liang et al. [29] develop performance models for inference on accelerators (e.g., GPU, TPU) and use them to guide DNN placement via an online knapsack algorithm. In contrast, our models are applicable not only to inference workloads but also to other accelerator-driven applications as shown in Section IV-D. They jointly capture device hardware characteristics, network delays, and device-to-edge performance differences, enabling direct comparison between on-device processing and edge offloading.
Modeling distributed edge workloads. Several studies have developed analytic models for distributed device–edge–cloud environments. Raei et al. [59] analyze the impact of wireless signal strength on request rejection probability and mean response delay. FogQN [60] uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers. Loghin et al. [61] model MapReduce applications in hybrid edge–cloud environments. However, these works abstract away accelerator-specific queuing behavior and do not consider the effects of multi-tenancy at the edge. Our work complements existing efforts by addressing these missing dimensions through a unified queuing-based modeling framework that informs resource management decisions.
VII Conclusions
This paper presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators. We validated the accuracy of our model across a range of hardware, workloads, and network configurations. We developed a resource manager to demonstrate how model predictions can be used to dynamically adapt the execution strategy in response to changing network conditions and multi-tenant edge server workloads. As future work, we plan to extend the framework to model continuous batching in LLM inference and to support multi-stage pipelines or DAGs with dependencies across stages.

### v3

@@ 4.8. Impact of Multi-Tenancy
Figure 5c compares the average latency of the two strategies as the number of co-located InceptionV4 applications on the edge increases, where each application receives 2 RPS from its corresponding client. The result shows that when only a single application is deployed on the edge, offloading yields lower average latency than on-device processing. However, as the number of co-located applications increases, offloading latency increases as well. This increase is caused by resource contention among applications because the A2 GPU lacks isolation mechanisms. As a result, there is a crossover point at 18 applications, beyond which on-device processing becomes more efficient. Note that the on-device processing latency remains stable regardless of the number of concurrent applications, as each device operates independently. In all cases, our models closely predict the latencies for both strategies.
@@
Key takeaway. Offloading average latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
5. Models in Action
@@
In this section, we demonstrate how a resource manager can leverage our models to adaptively switch between offloading and on-device processing. We outline its algorithm and demonstrate its effectiveness through two case studies that showcase how the models enable adaptation to network variability and to multi-tenant edge servers.
5.1. Model-Driven Adaptive Resource Management
@@
1
1ex// Predict on-device processing latency
@@ 11  $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
The resource manager runs locally on the device and periodically collects runtime metrics including network bandwidth, edge server load, and request arrival rate as discussed in Sec 4.2. At each epoch, it inputs these metrics into our models to estimate the average latency for each strategy, and selects the strategy with the lowest predicted latency to execute requests. Algorithm 1 outlines the decision-making process of the resource manager. Lines 1-2 predict the average latency of on-device processing based on the current arrival rate and the device’s estimated service time. Lines 3-6 compute the predicted offloading latency for each edge server, considering server load, service rates, and network conditions. Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
5.2. Case 1: Fluctuating Network Conditions
@@
Figure 6. Latencies under varying network bandwidth as the resource manager selects between the two strategies.
@@
We first demonstrate how the resource manager can leverage our models to adapt to dynamic network conditions, using TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS. Figure 6 shows the observed latency using the execution strategies chosen by the resource manager, alongside the model-predicted latency of each strategy. Initially, the device is configured with a 20 Mbps network to represent a stable 5G connection. Since the predicted offloading latency is lower than on-device processing, the manager chooses to offload requests. To emulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation. Here, despite the reduced bandwidth from 20 to 10 Mbps, the predicted latency for offloading remains lower, and the execution strategy remains unchanged. At t = 40, we further reduce the bandwidth to 2 Mbps to emulate poor coverage. Since the predicted offloading latency (45 ms) now exceeds the predicted on-device processing latency (21 ms), the resource manager switches to on-device processing. Finally, we restore the bandwidth to 20 Mbps at t = 60 to represent signal recovery, and the resource manager switches back to offloading. Throughout this sequence, our models closely predict the observed latencies, validating their accuracy.
@@
Key takeaway. Our models enable dynamic adaptation to changing network conditions by switching between execution strategies.
5.3. Case 2: Multi-tenant Edge Servers
Figure 7. Latencies under varying edge server workloads as the resource manager adapts.
@@ 5.3. Case 2: Multi-tenant Edge Servers
We now showcase how the resource manager can leverage our models to adapt to workload changes across multi-tenant edge servers. We use two edge servers, $E_{1}$ and $E_{2}$, each serving YOLOv8n inference requests using an A2 GPU. Initially, the server request rates are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$. Figure 7 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing. At t = 0, since the predicted latency for offloading to $E_{1}$ is the lowest, the resource manager offloads requests to $E_{1}$. At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50 to emulate workload fluctuations. In response, the resource manager adapts by offloading to the less loaded server $E_{2}$. Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50. Now, as the predicted latency of on-device processing becomes the lowest when both servers have a request rate of 50 RPS, the resource manager shifts to on-device processing. This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workloads.
@@
6. Related Work
Edge offloading. Edge offloading is widely adopted in many applications to enable resource-constrained devices to perform compute-intensive tasks by leveraging edge resources (Davis2004; Jang2018; Lai2017; Li2024; Ogden2021; Ren2019; Mei2012; Liu2018; zhou2017; Wang2018; Zhang2024MILCOM; Tkachenko2023; Ogden2023; Zilic2025). For example, Furion (Lai2017) selectively offloads compute-intensive components of mobile AR pipelines to the edge. Li et al. (Li2024) introduce an evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints. In contrast, our work uses a model-driven approach to provide end-to-end latency predictions and quantitative performance crossover predictions for accelerator-driven workloads and generalizes to device-edge collaborative processing.
Model-aware resource allocation. Queuing models have been widely used to predict the performance of traditional CPU-bound workloads (Jiang2021; WangB2024; Tantawi1985; Ambati2020; Gandhi2019; Shen2011; Urgaonkar2005). More recently, there has been growing interest in modeling the performance of hardware accelerators. For example, Merck et al. (Merck2019) model GPU batching latency in DNN inference workloads. Liang et al. (Liang2023) develop performance models for inference on accelerators to guide DNN placement via an online knapsack algorithm. In contrast, our models apply not only to inference workloads but also to other accelerator-driven workloads as shown in Sec 4.4.
Modeling distributed workloads. Several studies have developed analytic models for distributed device–edge–cloud environments. FogQN (Tadakamalla2018) uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers. Loghin et al. (Loghin2019) model MapReduce applications in hybrid edge–cloud environments. However, they abstract away accelerator-specific queuing behavior and do not account for multi-tenancy at the edge. Our work complements existing efforts by addressing these missing dimensions through a unified queuing-based modeling framework that informs resource management decisions.
7. Conclusions
This paper presented analytic models for on-device processing and edge offloading with accelerators, providing performance predictions for adaptive execution decisions. We validated their accuracy across a range of scenarios and developed a resource manager that applies these predictions to adapt to real-world dynamics. As future work, we plan to extend the models to support multi-stage pipelines represented as DAG structures, capturing inter-stage dependencies.